3. Two horns emit sounds of frequency 365 Hz and 371 Hz. What is their beat frequency?
4. An organ pipe manufacturer wants to make a pipe that resonates with a fundamental frequency of 32.75 Hz. Assume the speed of sound in air is 343 m/s. What must the length of the pipe be if it is
a. an open pipe?
c. a closed pipe?
1
Expert's answer
2020-04-23T12:58:46-0400

3.


"f_b=371-365=6\ Hz"


4.

a.


"l=\frac{v}{2f}=\frac{343}{2\cdot 32.75}=5.24\ m"

b.


"l=\frac{v}{4f}=\frac{343}{4\cdot 32.75}=2.62\ m"